A snowman stands on the ground. It consists of two solid spherical snow balls with radii R and r = R/2, respectively. The balls are placed directly on top of each other, with the smaller one on the top. The snow balls have the same constant density. At what height above the ground is the center of mass of the snowman located?
Answer #1

Let the mass of the smaller sphere be m , then the mass of the bigger sphere will be 8*m .

Center of mass H = (m1*h1+m2h2)/[m1+m2]

= (m*2.5R+8m*R)/[m+8m]

= 10.5mR/9m

= 10.5/9 R

= 1.12R